OpenStudy (anonymous):

find the solutions to the system. y=x^2+5x+6 y=4x+12

OpenStudy (babalooo):

4x+12 = x^2+5x+6 0 = x^2 +x -6 Factor and solve for X.

OpenStudy (anonymous):

x= -3,2 ??

OpenStudy (anonymous):

@babalooo

OpenStudy (anonymous):

Pls see attached.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Welcome. Verified with wolfram-alpha. http://www.wolframalpha.com/input/?i=solve+y+%3D+x%5E2+%2B+5x+%2B6++and+y%3D4x%2B12

OpenStudy (anonymous):

I have one more. y=x^2+8x+2 y=7x+4

OpenStudy (anonymous):

would that be no solutions??

OpenStudy (anonymous):

what website is that?

OpenStudy (anonymous):

http://www.wolframalpha.com/ Many commands available. Plot, solve good for you at present.

OpenStudy (babalooo):

Yup it's exactly what @BillBell posted. Just equate both equation since both = y. and factor for x. The put together equation would be \[x^{2} +x -2\]

OpenStudy (babalooo):

Wolframalpha is a good tool for math computation, but it's not something you should rely too much on... especially when it's just factoring... it takes away the practice/work that you're supposed to do.. it's a good tool for reference and graphs nonetheless.

OpenStudy (anonymous):

when i went to that website and typed the equation in it came up with a really long answer.

OpenStudy (anonymous):

@babaloo, you're a party pooper. Seriously though, when you're a student you need to learn how to do all of these things. But I'm retired and even when I was a graduate student (in mathematics) my arithmetic and algebra was dreadful. So I'm glad to have software to make sure I've done the manipulations correctly.

OpenStudy (anonymous):

wolfram-alpha tries to give you everything it thinks you might want. sometimes it doesn't give you anything you wanted though.

OpenStudy (babalooo):

Yes, I agree with you. I am also glad that there's these kinds of programs and software that help with computation, but if you consider the user who posted these factor problems, relying on these programs may become a habit and it wont help them in the long run for such fundamental math skills. I believe you have to pay to get long step by step work on wolframalpha.